OpenStudy (anonymous):

Calculate the de Broglie wavelength of a 143 g baseball traveling at 95 mph.

OpenStudy (paxpolaris):

\[\Large p=\frac h \lambda\]\[\therefore \lambda = \frac h {m \cdot v}\] http://lmgtfy.com/?q=h%2F(143g*95mph)